OpenStudy (anonymous):

sin2x+cos2x if tan=2 and cos> 0

OpenStudy (anonymous):

solve for cos with tan(squared) + 1 = sec(squared)

OpenStudy (anonymous):

4 + 1 = sec(squared) squareroot(5) = sec sec = 1/ cos 1/ squareroot(5) = cos

OpenStudy (anonymous):

now solve for sin with 1 + cot(squared) = csc(squared)

OpenStudy (anonymous):

1 + (1/4) = csc(squared) 5/4 = cc(squared) csc = squareroot(5)/2 csc = 1/ sin 2/ squareroot(5) = sin

OpenStudy (anonymous):

so sin 2x = 2 * sinx * cosx and cos 2x = cosx(squared) - sinx(squared)

OpenStudy (anonymous):

plug it in \[2 \times sin(2\div \sqrt{5}) \times cos(1\div \sqrt{5}) = \sin 2x\] \[\cos(1\div \sqrt{5})^{2} - \sin(2\div \sqrt{5}) = \cos 2x\]

OpenStudy (anonymous):

sin2x = 0.8 cos2x = -0.6 sin2x + cos2x 0.8 - 0.6 = 0.2